# On Edge Coloring of Multigraphs

### v4

@@ Abstract
Let $\Delta(G)$ and $\chi^{\prime}(G)$ be the maximum degree and chromatic index of a graph $G$, respectively. Appearing in different format, Gupta (1967), Goldberg (1973), Andersen (1977), and Seymour (1979) made the following conjecture: Every multigraph $G$ satisfies $\chi^{\prime}(G)\leq\max\{\Delta(G)+1,\Gamma(G)\}$, where $\Gamma(G)=\max_{H\subseteq G}\left\lceil\frac{|E(H)|}{\lfloor\tfrac{1}{2}|V(H)|\rfloor}\right\rceil$ is the density of $G$. In this paper, we present a polynomial-time algorithm for coloring any multigraph with $\max\{\Delta(G)+1,\Gamma(G)\}$ many colors, confirming the conjecture algorithmically. Since $\chi^{\prime}(G)\geq\max\{\Delta(G),\Gamma(G)\}$, this algorithm gives a proper edge coloring that uses at most one more color than the optimum. As determining the chromatic index of an arbitrary graph is $NP$-hard, the $\max\{\Delta(G)+1,\Gamma(G)\}$ bound is best possible for efficient proper edge coloring algorithms on general multigraphs, unless $P=NP$.
@@ 1 Introduction
A (proper) $k$-edge-coloring of $G$ is an assignment of $k$ colors, $1,2,\ldots,k$, to the edges of $G$ so that no two adjacent edges have the same color. By definition, the chromatic index $\chi^{\prime}(G)$ of $G$ is the minimum $k$ for which $G$ has a $k$-edge-coloring. We use $[k]$ to denote the color set $\{1,2,\ldots,k\}$, and use ${\cal C}^{k}(G)$ to denote the set of all $k$-edge-colorings of $G$. Note that every $k$-edge-coloring of $G$ is a mapping from $E$ to $[k]$. Let $\Delta$ and $\mu$ be the maximum degree and the multiplicity of a graph $G$, respectively. Vizing Vizing64 showed that $\Delta\leq\chi^{\prime}\leq\Delta+\mu$. Simple graphs are consequently divided into two classes. Class I: graphs with $\chi^{\prime}=\Delta$; Class II: graphs with $\chi^{\prime}=\Delta+1$.
Since the edges of $G$ with the same color form a matching, we have $|E(H)|\leq\chi^{\prime}(G)\lfloor|V(H)|/2\rfloor$ for any subgraph $H$ of $G$. Therefore, we have $\chi^{\prime}(G)\geq\max\{\Delta(G),\Gamma(G)\}$, where
$\Gamma(G):=\max_{H\subseteq G}\left\lceil\frac{|E(H)|}{\lfloor\frac{1}{2}|V(H)|\rfloor}\right\rceil,$
@@
On the other hand, Gupta  (1967) Gupta67 , Goldberg (1973) Goldberg , Andersen (1977) Andersen and Seymour (1979) Seymour independently made the following conjecture, commonly known as the Goldberg-Seymour conjecture:
@@
Over the past five decades, this conjecture has been a subject of extensive research in the fields of graph theory, theoretic computer science and operations research, and has inspired a significant body of work, with contributions from many researchers; see StiebSTF-Book by Stiebitz et al. and gs by Chen, Zang, and the author for a comprehensive account. In particular, gs confirms the Goldberg-Seymour conjecture. However, the approach used in gs is quite technical and artificial, and thus it is far from obtaining an efficient edge coloring algorithm. In fact, the time complexity is exponential on the number of vertices, which makes it difficult to deploy gs in real world applications.
In this paper, we confirm Conjecture 1 in a natural approach by finding a polynomial-time algorithm to properly edge color any (multi)graph $G$ with $\max\{\Delta(G)+1,\Gamma(G)\}$ many colors. Since $\chi^{\prime}(G)\geq\max\{\Delta(G),\Gamma(G)\}$, it is equivalent to saying that:
@@
Holyer Holyer81 proved that it is $NP$-hard in general to determine $\chi^{\prime}(G)$, even when restricted to a simple cubic graph. In this sense, the $\max\{\Delta(G)+1,\Gamma(G)\}$ bound is best possible for efficient edge coloring algorithms. Theorem 1.1 also confirms the following conjecture by Hochbaum, Nishizeki, and Shmoys HochNS80 .
@@ 2 Notation and terminology
We will generally follow StiebSTF-Book ; gs for notation and terminology. Let $G=(V,E)$ be a multigraph. For each $X\subseteq V$, let $G[X]$ denote the subgraph of $G$ induced by $X$, and let $G-X$ denote $G[V-X]$; we write $G-x$ for $G-X$ if $X=\{x\}$. Moreover, we use $\partial(X)$ to denote the set of all edges with precisely one end in $X$, and write $\partial(x)$ for $\partial(X)$ if $X=\{x\}$. Edges in $\partial(X)$ are called boundary edges of $X$. For each pair $x,y\in V$, let $E(x,y)$ denote the set of all edges between $x$ and $y$. As it is no longer appropriate to represent an edge $f$ between $x$ and $y$ by $xy$ in a multigraph, we write $f\in E(x,y)$ instead. For each subgraph $H$ of $G$, let $V(H)$ and $E(H)$ denote the vertex set and edge set of $H$, respectively, let $|H|=|V(H)|$, and let $G[H]=G[V(H)]$ and $\partial(H)=\partial(V(H))$.
@@
Let $e$ be an edge of $G$. A tree-sequence with respect to $G$ and $e$ is a sequence $T=(y_{0},e_{1},y_{1},\ldots,e_{p},y_{p})$ with $p\geq 1$, consisting of distinct edges $e_{1},e_{2},\ldots,e_{p}$ and distinct vertices $y_{0},y_{1},\ldots,y_{p}$, such that $e_{1}=e$ and each edge $e_{j}$ with $1\leq j\leq p$ is between $y_{j}$ and some $y_{i}$ with $0\leq i<j$. So a tree-sequence $T$ indeed induces a tree as a subgraph of $G$. Given a tree-sequence $T=(y_{0},e_{1},y_{1},\ldots,e_{p},y_{p})$, we can naturally associate a linear order $\prec$ with its vertices, such that $y_{i}\prec y_{j}$ if $i<j$. We write $y_{i}\preceq y_{j}$ if $i\leq j$. This linear order will be used repeatedly in subsequent sections. For each vertex $y_{j}$ of $T$ with $j\geq 1$, let $T(y_{j})$ denote $(y_{0},e_{1},y_{1},\ldots,e_{j},y_{j})$. Clearly, $T(y_{j})$ is also a tree-sequence with respect to $G$ and $e$. We call $T(y_{j})$ the segment of $T$ induced by $y_{j}$. Let $T_{1}$ and $T_{2}$ be two tree-sequences with respect to $G$ and $e$. We write $T_{1}\subseteq T_{2}$ if $T_{1}$ is a segment of $T_{2}$, and write $T_{1}\subset T_{2}$ if $T_{1}$ is a proper segment of $T_{2}$; that is, $T_{1}\subseteq T_{2}$ and $T_{1}\neq T_{2}$.
@@
Let $\varphi$ be a $k$-edge-coloring of $G$. For each $\alpha\in[k]$, the edge set $E_{\varphi,\alpha}=\{e\in E:\,\varphi(e)=\alpha\}$ is called a color class, which is a matching in $G$. For any two distinct colors $\alpha$ and $\beta$ in $[k]$, let $H$ be the spanning subgraph of $G$ with $E(H)=E_{\varphi,\alpha}\cup E_{\varphi,\beta}$. Then each component of $H$ is either a path or an even cycle; we refer to such a component as an $(\alpha,\beta)$-chain with respect to $\varphi$, and also call it an $(\alpha,\beta)$-path (resp. $(\alpha,\beta)$-cycle) if it is a path (resp. cycle). Possibly a component of $H$ is an isolated vertex. We use $P_{v}(\alpha,\beta,\varphi)$ to denote the unique $(\alpha,\beta)$-chain containing the vertex $v$. Clearly, for any two distinct vertices $u$ and $v$, $P_{u}(\alpha,\beta,\varphi)$ and $P_{v}(\alpha,\beta,\varphi)$ are either identical or vertex-disjoint. In this paper, $P_{v}(\alpha,\beta,\varphi)$ usually is a $(\alpha,\beta)$-path with $v$ as an end, unless specified otherwise. Let $C$ be an $(\alpha,\beta)$-chain with respect to $\varphi$, and let $\varphi^{\prime}$ be the $k$-edge-coloring arising from $\varphi$ by interchanging $\alpha$ and $\beta$ on $C$. We say that $\varphi^{\prime}$ is obtained from $\varphi$ by recoloring $C$, and write $\varphi^{\prime}=\varphi/C$. This operation is called a Kempe change.
@@
Let $F$ be an edge subset of $G$. As usual, $G-F$ stands for the multigraph obtained from $G$ by deleting all edges in $F$; we write $G-f$ for $G-F$ if $F=\{f\}$. Let $\varphi\in{\cal C}^{k}(G-F)$. For each $K\subseteq E$, define $\varphi(K)=\cup_{e\in K-F}\,\varphi(e)$. For each $v\in V$, define
$\varphi(v)=\varphi(\partial(v))\hskip 5.69054pt{\rm and}\hskip 5.69054pt\overline{\varphi}(v)=[k]-\varphi(v).$
@@
We call $X$ elementary with respect to $\varphi$ if $\overline{\varphi}(u)\cap\overline{\varphi}(v)=\emptyset$ for any two distinct vertices $u,v\in X$. We call $X$ closed with respect to $\varphi$ if $\varphi(\partial(X))\cap\overline{\varphi}(X)=\emptyset$; that is, no missing color of $X$ appears on the edges in $\partial(X)$. Furthermore, we call $X$ strongly closed with respect to $\varphi$ if $X$ is closed with respect to $\varphi$ and $\varphi(e)\neq\varphi(f)$ for any two distinct colored edges $e,f\in\partial(X)$. For each subgraph or tree-sequence $H$ of $G$, write $\overline{\varphi}(H)$ for $\overline{\varphi}(V(H))$, and write ${\varphi}(H)$ for ${\varphi}(E(H))$. If more than one boundary edges of $H$ (edges in $\partial(H)$) colored by $\alpha$, we call $\alpha$ a defective color of $H$ with respect to $\varphi$, call each $\alpha$-colored boundary edge a defective edge of $H$ with respect to $\varphi$, and call the end vertex of a defective edge in $H$ a defective vertex of $H$ with respect to $\varphi$. A color $\alpha\in\overline{\varphi}(H)$ is called closed in $H$ under $\varphi$ if $H$ has no $\alpha$-colored boundary edges. For convenience, we say that $H$ is closed (resp. strongly closed) with respect to $\varphi$ if $V(H)$ is closed (resp. strongly closed) with respect to $\varphi$. Let $\alpha$ and $\beta$ be two colors that are not assigned to $\partial(H)$ under $\varphi$. We use $\varphi/(G-H,\alpha,\beta)$ to denote the coloring $\varphi^{\prime}$ obtained from $\varphi$ by interchanging $\alpha$ and $\beta$ in $G-V(H)$. Since $\varphi$ belongs to ${\cal C}^{k}(G-F)$, so does $\varphi^{\prime}$.

### v5

@@ Abstract
Let $\Delta(G)$ and $\chi^{\prime}(G)$ be the maximum degree and chromatic index of a graph $G$, respectively. Appearing in different forms, Gupta (1967), Goldberg (1973), Andersen (1977), and Seymour (1979) made the following conjecture: Every multigraph $G$ satisfies $\chi^{\prime}(G)\leq\max\{\Delta(G)+1,\Gamma(G)\}$, where $\Gamma(G)=\max_{H\subseteq G,|V(H)|\geq 2}\left\lceil\frac{|E(H)|}{\lfloor\tfrac{1}{2}|V(H)|\rfloor}\right\rceil$ is the density of $G$. In this paper, we present a polynomial-time algorithm for coloring any multigraph with $\max\{\Delta(G)+1,\Gamma(G)\}$ colors, confirming the conjecture algorithmically. Since $\chi^{\prime}(G)\geq\max\{\Delta(G),\Gamma(G)\}$, this algorithm gives a proper edge coloring that uses at most one more color than the optimum. As determining the chromatic index of an arbitrary graph is $NP$-hard, the $\max\{\Delta(G)+1,\Gamma(G)\}$ bound is best possible for efficient proper edge coloring algorithms on general multigraphs, unless $P=NP$. Related work of Chen, Hao, Yu, and Zang have also obtained an algorithm using similar high-level ideas; the present approach establishes a complete proof.
@@ 1 Introduction
A (proper) $k$-edge-coloring of $G$ is an assignment of $k$ colors, $1,2,\ldots,k$, to the edges of $G$ so that no two adjacent edges have the same color. By definition, the chromatic index $\chi^{\prime}(G)$ of $G$ is the minimum $k$ for which $G$ has a $k$-edge-coloring. We use $[k]$ to denote the color set $\{1,2,\ldots,k\}$, and use ${\cal C}^{k}(G)$ to denote the set of all $k$-edge-colorings of $G$. Note that every $k$-edge-coloring of $G$ is a mapping from $E$ to $[k]$. Let $\Delta(G)$ and $\mu(G)$ be the maximum degree and the multiplicity of a graph $G$, respectively. Vizing [15] showed that $\Delta(G)\leq\chi^{\prime}(G)\leq\Delta(G)+\mu(G)$. Simple graphs are consequently divided into two classes. Class I: graphs $G$ with $\chi^{\prime}(G)=\Delta(G)$; Class II: graphs $G$ with $\chi^{\prime}(G)=\Delta(G)+1$.
Since the edges of $G$ with the same color form a matching, we have $|E(H)|\leq\chi^{\prime}(G)\lfloor|V(H)|/2\rfloor$ for any subgraph $H$ of $G$ with $|V(H)|\geq 2$. Therefore, we have $\chi^{\prime}(G)\geq\max\{\Delta(G),\Gamma(G)\}$, where
$\Gamma(G):=\max_{H\subseteq G,|V(H)|\geq 2}\left\lceil\frac{|E(H)|}{\lfloor\frac{1}{2}|V(H)|\rfloor}\right\rceil,$
@@
On the other hand, Gupta  (1967) [8], Goldberg (1973) [7], Andersen (1977) [1] and Seymour (1979) [12] independently made the following conjecture, commonly known as the Goldberg-Seymour conjecture:
@@
Over the past five decades, this conjecture has been a subject of extensive research in the fields of graph theory, theoretic computer science and operations research, and has inspired a significant body of work, with contributions from many researchers; see [13] by Stiebitz et al. and [6] by Chen, Zang, and the author for a comprehensive account. In particular, [6] confirms the Goldberg-Seymour conjecture. However, the approach used in [6] is quite technical and artificial, and thus it is far from obtaining an efficient edge coloring algorithm. In fact, the time complexity is exponential on the number of vertices, which makes it difficult to deploy [6] in real world applications.
The key to prove the Goldberg-Seymour conjecture is to find a “strongly closed elementary set”, see the proof of Theorem 3.3 for more details. The Tashkinov tree method, introduced by Tashkinov [14] in the early 2000s, is the perfect tool for this purpose. Therefore, almost all recent progress towards the conjecture is based on the extensions of Tashkinov trees. To properly extend a Tashkinov tree, two major questions need to be answered:
  1. (1)
    Where to extend. This often involves adding requirements on the tree sequence we extend from.
  2. (2)
    How to mimic Tashkinov’s idea to show that the extended structure is elementary. This involves (a) Kempe changes, and (b) proving structure and all the extension requirements stay well-defined after the Kempe changes.
Question (2) was answered naturally in [5] and [6] by utilizing the idea of interchangeability. This idea shows that under certain conditions, for a color $\alpha$ missing at a vertex inside an extended Tashkinov tree $T_{n}$ (that is a Tashkinov-type tree sequence extended n1n-1 times) and a color $\delta$ appearing on the boundary of $T_{n}$, there is at most one $(\alpha,\delta)$-path sharing vertices with $T_{n}$. All the other $(\alpha,\delta)$-components sharing vertices with $T_{n}$ must be cycles. In this way, we can mimic Tashkinov’s idea on proving the elementariness after extending $T_{n}$ by avoiding this unique $(\alpha,\delta)$-path sharing vertices with $T_{n}$.
On the other hand, Question (1) was not fully answered in [6]. Pick an arbitrary color $\alpha$ missing at an arbitrary vertex inside $T_{n}$ and a ”defective color” $\delta$ (defective means that there are more than one edge colored by $\delta$ that appears on the boundary of $T_{n}$). Ideally, we should be able to naturally extend $T_{n}$ with any $\delta$-colored defective boundary edges except for the one that is incident with the $(\alpha,\delta)$-exit vertex of $T_{n}$, which is the last vertex of the only $(\alpha,\delta)$-path (uniqueness from interchangeability) shared with $T_{n}$ along the direction starting inside $T_{n}$. One can refer to the later Figure 1 for a graphical explanation. Asplund and McDonald [2] also suggested a similar type of extension. However, [6] cannot guarantee that the extension structure is well-defined after performing Kempe changes. To be more specific, after Kempe changes involving the color $\alpha$ or $\delta$, the $(\alpha,\delta)$-path intersecting $T_{n}$ might change its exit. So the entire proof collapses. To overcome this obstacle, [6] introduced a sophisticated way of extending a Tashkinov tree with a lot of “maximum” rules while changing the coloring of $G$ along the extension process. This does work for the proof, but it is impossible to find a corresponding edge coloring of $G$ in polynomial time.
To address this problem, in this paper we showed that either such Kempe changes do not change the exit of the aforementioned $(\alpha,\delta)$-path intersecting $T_{n}$, or there is a way to reduce the “size” (not the size of the vertex set) of a counterexample. See Lemmas 4.1 to 4.7 and their proofs for more details on this. So Question (1) is answered naturally, which leads to a polynomial-time edge coloring algorithm.
In particular, we confirm Conjecture 1 by finding a polynomial-time algorithm to properly edge color any (multi)graph $G$ with $\max\{\Delta(G)+1,\Gamma(G)\}$ many colors. Since $\chi^{\prime}(G)\geq\max\{\Delta(G),\Gamma(G)\}$, it is equivalent to saying that:
@@
Holyer [11] proved that it is $NP$-hard in general to determine $\chi^{\prime}(G)$, even when restricted to a simple cubic graph. In this sense, the $\max\{\Delta(G)+1,\Gamma(G)\}$ bound is best possible for efficient edge coloring algorithms. Theorem 1.1 also confirms the following conjecture by Hochbaum, Nishizeki, and Shmoys [10].
@@
It is worth noting that in July 2024 (the first draft of this paper was posted on arXiv in August 2023), Chen, Hao, Yu, and Zang [4] also posted a related result on arXiv using a similar approach. In their draft, Lemmas 4.34.7 are combined into a single lemma without incorporating the techniques developed in Lemma 4.6. As explained in Remark 2, this omission leaves a critical case unresolved.
@@ 2 Notation and terminology
We will generally follow [13, 6] for notation and terminology. Let $G=(V,E)$ be a multigraph. For each $X\subseteq V$, let $G[X]$ denote the subgraph of $G$ induced by $X$, and let GXG-X denote G[VX]G[V-X]; we write GxG-x for $G-\{x\}$. Moreover, we use $\partial(X)$ to denote the set of all edges with precisely one end in $X$, and write $\partial(x)$ for $\partial(\{x\})$. Edges in $\partial(X)$ are called boundary edges of $X$. For each pair $x,y\in V$, let $E(x,y)$ denote the set of all edges between $x$ and $y$. As it is no longer appropriate to represent an edge $f$ between $x$ and $y$ by $xy$ in a multigraph, we write $f\in E(x,y)$ instead. For each subgraph $H$ of $G$, let $V(H)$ and $E(H)$ denote the vertex set and edge set of $H$, respectively, let $|H|=|V(H)|$, and let $G[H]=G[V(H)]$ and $\partial(H)=\partial(V(H))$.
@@
Let $e$ be an edge of $G$. A tree-sequence with respect to $G$ and $e$ is a sequence $T=(y_{0},e_{1},y_{1},\ldots,e_{p},y_{p})$ with $p\geq 1$, consisting of distinct edges $e_{1},e_{2},\ldots,e_{p}$ and distinct vertices $y_{0},y_{1},\ldots,y_{p}$, such that $e_{1}=e$ and each edge $e_{j}$ with $1\leq j\leq p$ is between $y_{j}$ and some $y_{i}$ with $0\leq i<j$. Let $E(T)$ and $V(T)$ be the edge set and vertex set of a tree-sequence $T$, respectively. So $E(T)$ indeed induces a tree as a subgraph of $G$. Given a tree-sequence $T=(y_{0},e_{1},y_{1},\ldots,e_{p},y_{p})$, we can naturally associate a linear order $\prec$ with its vertices, such that $y_{i}\prec y_{j}$ if $i<j$. We write $y_{i}\preceq y_{j}$ if $i\leq j$. This linear order will be used repeatedly in subsequent sections. For each vertex $y_{j}$ of $T$ with $j\geq 1$, let $T(y_{j})$ denote $(y_{0},e_{1},y_{1},\ldots,e_{j},y_{j})$. Clearly, $T(y_{j})$ is also a tree-sequence with respect to $G$ and $e$. We call $T(y_{j})$ the segment of $T$ induced by $y_{j}$. Let $T_{1}$ and $T_{2}$ be two tree-sequences with respect to $G$ and $e$. We write $T_{1}\subseteq T_{2}$ if $T_{1}$ is a segment of $T_{2}$, and write $T_{1}\subset T_{2}$ if $T_{1}$ is a proper segment of $T_{2}$; that is, $T_{1}\subseteq T_{2}$ and $T_{1}\neq T_{2}$.
@@
Let $\varphi$ be a $k$-edge-coloring of $G$. For each $\alpha\in[k]$, the edge set $E_{\varphi,\alpha}=\{e\in E:\,\varphi(e)=\alpha\}$ is called a color class, which is a matching in $G$ if it is nonempty. For any two distinct colors $\alpha$ and $\beta$ in $[k]$, let $H$ be the spanning subgraph of $G$ with $E(H)=E_{\varphi,\alpha}\cup E_{\varphi,\beta}$. Then each component of $H$ is either a path or an even cycle; we refer to such a component as an $(\alpha,\beta)$-chain with respect to $\varphi$, and also call it an $(\alpha,\beta)$-path (resp. $(\alpha,\beta)$-cycle) if it is a path (resp. cycle). Possibly a component of $H$ is an isolated vertex. We use $P_{v}(\alpha,\beta,\varphi)$ to denote the unique $(\alpha,\beta)$-chain containing the vertex $v$. Clearly, for any two distinct vertices $u$ and $v$, $P_{u}(\alpha,\beta,\varphi)$ and $P_{v}(\alpha,\beta,\varphi)$ are either identical or vertex-disjoint. In this paper, $P_{v}(\alpha,\beta,\varphi)$ usually is an $(\alpha,\beta)$-path with $v$ as an end, unless specified otherwise. Let $C$ be an $(\alpha,\beta)$-chain with respect to $\varphi$, and let $\varphi^{\prime}$ be the $k$-edge-coloring arising from $\varphi$ by interchanging $\alpha$ and $\beta$ on $C$. We say that $\varphi^{\prime}$ is obtained from $\varphi$ by recoloring $C$, and write $\varphi^{\prime}=\varphi/C$. This operation is called a Kempe change.
@@
Let $F$ be an edge subset of $G$. As usual, GFG-F stands for the multigraph obtained from $G$ by deleting all edges in $F$; we write GfG-f for $G-\{f\}$. Let $\varphi\in{\cal C}^{k}(G-F)$. For each $K\subseteq E$, define $\varphi(K)=\cup_{e\in K-F}\,\varphi(e)$. For each $v\in V$, define
$\varphi(v)=\varphi(\partial(v))\quad{\rm and}\quad\overline{\varphi}(v)=[k]-\varphi(v).$
@@
We call $X$ elementary with respect to $\varphi$ if $\overline{\varphi}(u)\cap\overline{\varphi}(v)=\emptyset$ for any two distinct vertices $u,v\in X$. We call $X$ closed with respect to $\varphi$ if $\varphi(\partial(X))\cap\overline{\varphi}(X)=\emptyset$; that is, no missing color of $X$ appears on the edges in $\partial(X)$. Furthermore, we call $X$ strongly closed with respect to $\varphi$ if $X$ is closed with respect to $\varphi$ and $\varphi(e)\neq\varphi(f)$ for any two colored edges $e,f\in\partial(X)$ with $e\neq f$. For each subgraph or tree-sequence $H$ of $G$, write $\overline{\varphi}(H)$ for $\overline{\varphi}(V(H))$, and write ${\varphi}(H)$ for ${\varphi}(E(H))$. If more than one boundary edges of $H$ (edges in $\partial(H)$) colored by $\alpha$, we call $\alpha$ a defective color of $H$ with respect to $\varphi$, call each $\alpha$ colored boundary edge a defective edge of $H$ with respect to $\varphi$, and call the end vertex of a defective edge in $H$ a defective vertex of $H$ with respect to $\varphi$. A color $\alpha\in\overline{\varphi}(H)$ is called closed in $H$ under $\varphi$ if $H$ has no $\alpha$ colored boundary edges. For convenience, we say that $H$ is closed (resp. strongly closed) with respect to $\varphi$ if $V(H)$ is closed (resp. strongly closed) with respect to $\varphi$. Let $\alpha$ and $\beta$ be two colors that are not assigned to $\partial(H)$ under $\varphi$. We use $\varphi/(G-H,\alpha,\beta)$ to denote the coloring $\varphi^{\prime}$ obtained from $\varphi$ by interchanging $\alpha$ and $\beta$ in GV(H)G-V(H). Since $\varphi$ belongs to ${\cal C}^{k}(G-F)$, so does $\varphi^{\prime}$.